OpenStudy (anonymous):

Help for a medal?!?! Simplify. 5 square root of 11 end root minus 12 square root of 11 end root minus 2 square root of 11

OpenStudy (anonymous):

\[5\sqrt{11}-12\sqrt{11} - 2\sqrt{11} = (5-12-2 )\sqrt{11}\] You can factor out the sqrt(11)

OpenStudy (anonymous):

= -9 square root of 11

OpenStudy (anonymous):

Sorry for the delay. But thank you again for the help!